# A Local Projection Stabilised HHO Method for the Oseen Problem

### v1

@@ Abstract
Fluid flow problems with high Reynolds number show spurious oscillations in their solution when solved using standard Galerkin finite element methods. These Oscillations can be eradicated using various stabilisation techniques. In this article, we use a local projection stabilisation for a Hybrid High-Order approximation of the Oseen problem. We prove an existence-uniqueness result under a SUPG-like norm. We derive an optimal order error estimate under this norm for equal order polynomial discretisation of velocity and pressure spaces.
@@ Introduction
where $\boldsymbol{u}$ denotes the velocity of the fluid and $p$ denotes the pressure. Here $\Omega\subset\mathbb{R}^{2}$ is a polygonal domain with boundary $\partial\Omega$. The force function $\boldsymbol{f}$ is in $[L^{2}(\Omega)]^{2}$. The viscosity coefficient is denoted by $\epsilon$, where $0<\epsilon\ll 1$. The convection coefficient $\boldsymbol{b}$ is a $[W^{1}_{\infty}(\Omega)]^{2}$ function such that ${\rm div~{}}\boldsymbol{b}=0$. The reaction coefficient is a positive constant denoted by $\sigma$ with $\sigma\leq\sigma_{0}$.
@@
Fluid flow problems with dominant convection produce boundary and interior layers. It is well-known that the numerical solution to these problems using the usual Galerkin method cannot capture these small layers. Instead, they produce nonphysical solutions which contain spurious oscillations. To eliminate the effect of convection, one can add stabilisations. In this article, we focus on local projection stabilisation using a Hybrid High-Order approximation on general polygonal meshes.
@@
In the last few years, there has been a growing interest in high-order polynomial approximations of solutions to PDEs on general polytopal meshes. Due to the vast literature in this area, we cite a few well-known works: the Hybridizable Discontinuous Galerkin (HDG) method in Cock_Gopal_Laz_09_Unif ; Piet_Ern_12_DG_book . The HDG has been further extended to the convection diffusion problem in Cock_Dong_09_HybridDG and the Oseen problem in Oseen_HDG_cock . The Virtual Element method (VEM) has been studied in Beira_Brezz_Cang_Man_Mar_13_VEM ; Beira_Brezz_Mari_13_VEM_Elast ; Brezz_Falk_Mari_14_VEM_Mixed . The VEM has also been applied to the convection diffusion problem with SUPG stabilisation VEM_con_diff_supg and to the Oseen problem with LPS stabilisation Oseen_VEM_LPS . The Weak Galerkin method is introduced in Mu_Wang_Ye_15_WeakGaler_Nonlin ; Wang_Ye_13_Weak_Galerkin ; Wang_Ye_14_Mixed_Weak_Galerkin and the Gradient Discretisation method is introduced in Dron_Eym_Herb_GDM_16 ; Dron_Eym_GDM_18_book ; Piet_Dron_Man_18_GradDis_Polytope . The Multiscale Hybrid-Mixed method has been studied in Arya_Hard_Pare_Vale_13_Multiscale . The focus of our article is on the Hybrid High-Order (HHO) method, originally introduced in Piet_Ern_Lem_14_arb_local ; ern_pietro_elasticity . For an overview of the HHO method, we refer to Piet_Jero_HHO_Book_20 . HHO is a robust method based on a local polynomial reconstruction. It is independent of the dimension of the problem and suitable for local static condensation, which drastically reduces the computational cost of the matrix solver.
@@
The HHO method is closely related to the HDG method but they differ in the choice of stabilisation, see Cock_Piet_Ern_16_HHO_HDG for details. In the nonconforming Virtual Element methods (ncVEM) one takes the projection of the virtual function in the stabilisation, whereas the HHO method takes a reconstruction of the function in the stabilisation. In Lem_21_Bridging_HHO_VEM , the connection of the HHO method with the virtual element method is discussed. See Bone_Ern_14_poly ; Dron_Eym_Gall_10_unified ; Brezzi_Lip_Shas_05_Conv_Mimetic ; Brezzi_Lip_Shas_07_polyhedral ; Kuzn_Lip_Shas_04_Mimetic_poly for related works. In the lowest-order case ($k=0$), the HHO method resembles the Hybrid Mixed Mimetic family, hence the mixed-hybrid Mimetic Finite Differences, the Hybrid Finite Volume and the Mixed Finite Volume methods too, see Dron_Eym_Gall_Her_10_Mimetic ; Brez_Lipn_Simo_05_Mim_Dif ; Eym_Gall_Herb_10_NonconfMesh ; Dron_14_FV_Diff ; Dron_Eym_06_MixedFV . We state the following predominant works on the HHO methods: problem with pure diffusion Piet_Ern_Lem_14_arb_local , advection-diffusion problem Piet_Dron_Ern_15_HHO_Adv_Diff_Rea , interface problems Burman_Ern_18_HHO_Interface , for linear PDEs, elliptic obstacle problem Cicu_Ern_Gudi_20_HHO_Obstacle , Stokes problem Piet_Ern_Link_16_HHO_Stokes , the Oseen problem Oseen_HHO and the steady incompressible Navier Stokes equations Piet_Krell_18_NSE .
@@
In the study of stabilisations for fluid flow problems with high Reynolds number, the SUPG method by Hughes and Brooks MR679322 is the most well-known. SUPG is studied for the incompressible Navier Stokes equation in MR1377246 . There is a wide range of stabilisation techniques in the literature, some of them are: the method of least squares MR2454024 , residual free bubbles technique MR2454024 , continuous interior penalty method MR2454024 and the discontinuous Galerkin method MR2454024 . In this article, we are interested in the local projection stabilisation scheme, originally introduced for the Stokes problem by Becker and Braack MR1890352 for the Stokes problem. It has also been studied for the transport problem, scalar convection diffusion problem, the Oseen problem and the Navier stokes equations MR2121360 ; MR2454024 ; BRAACK2006372 ; MR2206447 ; MR2242376 . A non-conforming patchwise LPS scheme using Crouzeix-Raviart elements for the convection diffusion problem has been studied in MR3915226 and for the Oseen problem in biswas2019edge . A generalised version of LPS has been studied by Knobloch for the convection diffusion problem MR2670000 and also for the Oseen problem MR3003108 .
@@
The SUPG method naturally gives an additional control on the advective derivative of velocity; however, the usual LPS schemes in MR2121360 ; MR2206447 do not provide this. Moreover, the LPS schemes in MR2121360 ; MR2206447 work through a two-level mesh approach or through enrichment. The articles MR2121360 ; MR2206447 ; MR2670000 ; MR3003108 need to satisfy a local inf-sup condition necessary for error analysis and stability. In this article, we employ a generalised LPS technique to design an HHO scheme for the Oseen problem motivated by the work in Oseen_HHO on the HHO approximation for the Oseen problem. The additional LPS term provides control on the advective derivative. Moreover, we employ a one-level approach that does not require any enrichment of the discrete spaces and the need for a local inf-sup condition as seen in MR3003108 .
@@
In this article, along with the LP stabilisation, we have added another velocity stabilisation to control the normal jump of the solution. This helps to further stabilize the solution. Moreover, we also need pressure stabilisation to stabilise the gradient of pressure. In a nutshell, the LPS-HHO scheme is a combination of a usual HHO method for the Oseen problem combined with the above stabilisations along with an upwind term. Comparing this article with Oseen_HHO , we have proven that the LPS stabilisation term in the formulation helps to prove a stability result under a stronger SUPG-like norm. Moreover, in this article, the presence of normal jump stabilisation in the discrete scheme gives epsilon robust error bounds. This can be seen in the inequality (4).
@@
The rest of the article is organised as follows: Section 1 defines the Oseen problem along with some notation and preliminaries. Section 2 deals with discrete HHO formulation of the Oseen equations. Section 3 provides the proof for the discrete well-posedness of the system in Section 2. Section 5 provides the a priori error estimates. Numerical results are provided in Section 6. From now on, we denote by $a\lesssim b$ the expression $a\leq Cb$, where $C$ is a positive constant. The analysis is done on standard $k$th order Sobolev spaces $H^{k}(\Omega)$. Sobolev spaces with zero trace is denoted by $H^{k}_{0}(\Omega)$ with the standard norm $\left\|\cdot\right\|_{k}=\sum_{|\alpha|\leq k}\left\|D^{\alpha}\cdot\right\|_{L^{2}(\Omega)}$. $L^{2}_{0}(\Omega)$ represents the space of square integrable functions with zero mean. For $M\subset\Omega,k\geq 0$ and $1\leq p\leq\infty$, let .$\|.\|_{k,p,M}$ be the norm on the $k-$th order Sobolev space $W^{k,p}(M)$. For $p=2$ we denote the norm by .$\|.\|_{k,M}$.
@@ 1 Continuous Problem, Notations and Preliminaries
In this section, we introduce the weak formulation of the Oseen problem (0.1) and some preliminaries. Let $\boldsymbol{V}=[H^{1}_{0}(\Omega)]^{2}$ be the velocity space and $Q=L^{2}_{0}(\Omega)$ be the pressure space. The weak formulation for the Oseen problem (0.1) is given by: Find $\boldsymbol{u}\in\boldsymbol{V}$ and $p\in Q$ such that
@@
$a(\boldsymbol{u},\boldsymbol{v})$ $:=\epsilon(\nabla\boldsymbol{u},\nabla\boldsymbol{v})_{L_{2}(\Omega)}+((\boldsymbol{b}\cdot\nabla)\boldsymbol{u},\boldsymbol{v})_{L_{2}(\Omega)}+(\sigma\boldsymbol{u},\boldsymbol{v})_{L_{2}(\Omega)},$
$b(q,\boldsymbol{u})$ $:=(q,{\rm div~{}}\boldsymbol{u})_{L_{2}(\Omega)}.$
Using the fact that ${\rm div~{}}\boldsymbol{b}=0$ and $\sigma>0$, one can show that the bilinear form $a(\boldsymbol{u},\boldsymbol{v})$ is coercive. It is well known that the bilinear form $b(q,\boldsymbol{u})$ is inf-sup stable for $\boldsymbol{u}\in[H^{1}_{0}(\Omega)]^{2}$ and $q\in L^{2}_{0}(\Omega)$. Therefore, the existence and uniqueness of the problem (1.1) can be shown using the Babuška-Brezzi condition, see (MR851383, , Chapter IV). An equivalent formulation for (1.1) seeks $(\boldsymbol{u},p)\in\boldsymbol{V}\times Q$ such that
@@
Consider a decomposition $\mathcal{T}_{h}$ of the domain $\Omega$ into a finite collection of nonempty, disjoint, open polygons. Let $h_{T}$ denote the diameter of a polygon $T\in\mathcal{T}_{h}$. The subscript $h$ in $\mathcal{T}_{h}$ denotes the maximum diameter among all polygons $T$, that is $h:=\max_{T\in\mathcal{T}_{h}}h_{T}$. The edges/faces of a polygon $T$ are denoted by $F$. The collection of all faces (skeleton ) of the decomposition $\mathcal{T}_{h}$ is denoted by $\mathcal{F}_{h}$. The set of all interior faces is denoted by $\mathcal{F}_{h}^{i}$ and the set of all boundary faces by $\mathcal{F}_{h}^{b}$. Length of a face $F$ is denoted by $h_{F}$. We assume that the diameter of the polygons in $\mathcal{T}_{h}$ are uniformly comparable to the face lengths, that is, $h_{T}\lesssim h_{F}\lesssim h_{T}$. For a polygon $T$, let the collection of edges of $T$ be denoted by $\mathcal{F}_{T}$. We assume that there exists a constant $N$ such that $\max_{T\in\mathcal{T}_{h}}\#\mathcal{F}_{T}\leq N$. This condition restricts the polygons from having too many faces.
Next, we define the hybrid discrete spaces on the decomposition $\mathcal{T}_{h}$. For any bounded domain $S$, let $P^{k}(S)$ denote the space of polynomials defined on $S$ of degree at most $k\geq 0$. The local degrees of freedom on each polygon $T\in\mathcal{T}_{h}$ is given by
$\underline{\boldsymbol{U}}_{T}^{k}:=\{\underline{\boldsymbol{v}}_{T}=(\boldsymbol{v}_{T},(\boldsymbol{v}_{F})_{F\in\mathcal{F}_{T}}):\boldsymbol{v}_{T}\in[P^{k}(T)]^{2}\text{ and }\boldsymbol{v}_{F}\in[P^{k}(F)]^{2},F\in\mathcal{F}_{T}\}.$
@@
The restriction of $\underline{\boldsymbol{v}}_{h}\in\underline{\boldsymbol{U}}_{T}^{k}$ on a polygon $T$ is denoted by $\boldsymbol{v}_{T}$. The local interpolation operator $\underline{\boldsymbol{I}}_{T}^{k}:[H^{1}(T)]^{2}\rightarrow\underline{\boldsymbol{U}}_{T}^{k}$ is given by
@@
where $\pi_{T}^{k}$ and $\pi_{F}^{k}$ are $L^{2}$ orthogonal projection onto $[P^{k}(T)]^{2}$ and $[P^{k}(F)]^{2}$ respectively. In a similar manner, the global interpolation operator $\underline{\boldsymbol{I}}_{h}^{k}\boldsymbol{v}:[H^{1}(\Omega)]^{2}\rightarrow\underline{\boldsymbol{U}}_{h}^{k}$ is defined as follows
@@
We recall some standard inequalities which will be used throughout the article.
Inverse Inequality: There exists a positive constant $C$ independent of the meshsize $h_{T}$ such that for any $v_{h}\in P^{k}(T)$ we have
@@
Trace Inequality:  (Piet_Ern_12_DG_book, , pp. 27) There exists a positive constant $C$ independent of the meshsize $h_{T}$ such that
@@
Discrete Poincare inequality: There exists a positive constant $C$ independent of $h_{T}$ such that for $\underline{\boldsymbol{v}}_{h}\in\underline{\boldsymbol{U}}_{h,0}^{k}$ we have
$\left\|\boldsymbol{v}_{h}\right\|\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}\left\|\nabla\boldsymbol{v}_{T}\right\|_{T}^{2}+\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\frac{1}{h_{F}}\int_{F}(\boldsymbol{v}_{F}-\boldsymbol{v}_{T})^{2}\Big{)}^{1/2}.$ (1.3)
Approximation property of $L^{2}$ orthogonal projection:  (Piet_Ern_12_DG_book, , lemma 1.58) The $L^{2}$-projection $\pi_{T}^{k}$ satisfies the following approximation property: for any $\boldsymbol{v}\in H^{s}(T)$ with $s\in\{1,2,...,k+1\}$
@@
Let the outward unit normal component for a polygon $T\in\mathcal{T}_{h}$ be denoted by $\boldsymbol{n}_{T}$. Similarly the outward unit normal for a face $F\in\mathcal{F}_{T}$ is given by $\boldsymbol{n}_{TF}$ such that $\boldsymbol{n}_{T}\rvert_{F}=\boldsymbol{n}_{TF}$. Moreover, the normal component of the convection term on a face $F\in\mathcal{F}_{T}$ is defined as $\boldsymbol{b}_{TF}:=\boldsymbol{b}\rvert_{F}\cdot\boldsymbol{n}_{TF}$. The jump of a scalar-valued function $v$ on a face $F$ shared by two polygons $T_{1}$ and $T_{2}$ is given by
@@
The sign of $\left[\hskip-3.5pt\left[v\right]\hskip-3.5pt\right]$ is adjusted according to the direction of the outward normal. For a vector-valued function $\boldsymbol{v}=(v_{1},v_{2})$
$\nabla\boldsymbol{v}:=\begin{pmatrix}\frac{\partial v_{1}}{\partial x}&\frac{\partial v_{1}}{\partial y}\\ \frac{\partial v_{2}}{\partial x}&\frac{\partial v_{2}}{\partial y}\end{pmatrix},\hskip 14.22636pt\frac{\partial\boldsymbol{v}}{\partial\boldsymbol{n}}:=\nabla\boldsymbol{v}\cdot\boldsymbol{n}\;\;{\rm and}\;\;(\nabla\boldsymbol{u},\nabla\boldsymbol{v})_{L^{2}(\Omega)}:=\int_{\Omega}\sum_{i,j=1}^{2}(\nabla\boldsymbol{u})_{i,j}(\nabla\boldsymbol{v})_{i,j}{\rm~{}dx}.$
Let $|v|$ denote the modulus function of $v$. The positive and negative part of $v$ is defined as

### v2

@@ Abstract
In this article, we consider a local projection stabilisation for a Hybrid High-Order (HHO) approximation of the Oseen problem. We prove an existence-uniqueness result under a stronger SUPG-like norm. We improve the stability and provide error estimation in stronger norm for convection dominated Oseen problem. We also derive an optimal order error estimate under the SUPG-like norm for equal-order polynomial discretisation of velocity and pressure spaces. Numerical experiments are performed to validate the theoretical results.
@@ Introduction
where $\boldsymbol{u}$ denotes the velocity of the fluid and $p$ denotes the pressure. Here $\Omega\subset\mathbb{R}^{d}$ with $d\in\{2,3\}$ is a bounded polytopal domain with Lipschitz boundary $\partial\Omega$. The force function $\boldsymbol{f}$ is in $[L^{2}(\Omega)]^{d}$. The viscosity coefficient is denoted by $\epsilon$, where $0<\epsilon\ll 1$. The convection coefficient $\boldsymbol{b}$ is a $[W^{1}_{\infty}(\Omega)]^{d}$ function such that ${\rm div~{}}\boldsymbol{b}=0$. The reaction coefficient is a positive constant denoted by $\sigma$.
@@
Fluid flow problems with dominant convection produce boundary and interior layers. It is well-known that the numerical solution to these problems using the usual Galerkin method cannot capture these small layers. Instead, they produce nonphysical solutions that contain spurious oscillations. To eliminate the effect of convection, one can add stabilisations. The Hybrid High-Order method with upwind stabilisation of [1] eradicates the oscillation. In this article, we focus on local projection stabilisation using a Hybrid High-Order approximation on general polygonal meshes and establish the stability and error estimation in a stronger norm.
@@
In the last few years, there has been a growing interest in high-order polynomial approximations of solutions to PDEs on general polytopal meshes. Due to the vast literature in this area, we cite a few well-known works: the Hybridizable Discontinuous Galerkin (HDG) method in [24, 28]. The HDG has been extended further to the convection diffusion problem in [23] and the Oseen problem in [19]. The Virtual Element Method (VEM) has been studied in [5, 6, 13]. The VEM has also been applied to the convection diffusion problem with SUPG stabilisation [7] and to the Oseen problem with LPS stabilisation [49]. The Weak Galerkin method is introduced in [50, 53, 54] and the Gradient Discretisation method is introduced in [39, 37, 27]. The Multiscale Hybrid-Mixed method has been studied in [2]. The focus of our article is on the Hybrid High-Order (HHO) method, originally introduced in [30, 29]. For an overview of the HHO method, we refer to [25]. HHO is a robust method based on local polynomial reconstruction. It is independent of the dimension of the problem and suitable for local static condensation, which drastically reduces the computational cost of the matrix solver.
@@
The HHO method is closely related to the HDG method, but differs in the choice of stabilisation; see [22] for details. In the nonconforming Virtual Element methods (ncVEM) one takes the projection of the virtual function in the stabilisation, whereas the HHO method takes a reconstruction of the function in the stabilisation. In [48], the connection of the HHO method with the virtual element method is discussed. See [9, 38, 14, 15, 47] for related works. In the lowest-order case ($k=0$), the HHO method resembles the Hybrid Mixed Mimetic family, hence the mixed-hybrid Mimetic Finite Differences, the Hybrid Finite Volume and the Mixed Finite Volume methods too, see [38, 16, 41, 35, 36]. We state the following predominant works on the HHO methods: problem with pure diffusion [30], advection-diffusion problem [26], interface problems [18], for linear PDEs, elliptic obstacle problem [21], Stokes problem [31], the Oseen problem [1], and the steady incompressible Navier-Stokes equations [32].
@@
In the study of stabilisations for fluid flow problems with a high Reynolds number, the SUPG method by Hughes and Brooks [17] is the most well known. SUPG is studied for the incompressible Navier-Stokes equations in [52]. There is a wide range of stabilisation techniques in the literature, some of them are: the least squares method [51], residual free bubbles technique [51], continuous interior penalty method [51], and the discontinuous Galerkin method [51]. In this article, we are interested in the local projection stabilisation scheme, originally introduced for the Stokes problem by Becker and Braack [3] for the Stokes problem. It has also been studied for the transport problem, the scalar convection diffusion problem, the Oseen problem, and the Navier-Stokes equations [4, 51, 11, 10, 12]. A nonconforming patchwise LPS method using Crouzeix-Raviart elements for the convection diffusion problem has been studied in [34] and for the Oseen problem in [8]. Knobloch has studied a generalised version of LPS for the convection diffusion problem in [45] and also for the Oseen problem [46].
@@
The SUPG method naturally gives an additional control over the advective derivative of the velocity; however, the usual LPS methods in [4, 10] do not provide this. Moreover, the LPS methods in [4, 10] work through a two-level mesh approach or through enrichment. The articles [4, 10, 45, 46] need to satisfy a local inf-sup condition necessary for error analysis and stability. In this article, we employ a generalised LPS technique to design an HHO method for the Oseen problem motivated by the works in [1] on the HHO approximation and in [46] on LPS stabilization for the Oseen problem. The additional LPS term provides control on the advective derivative. Moreover, we employ a one-level approach that does not require any enrichment of discrete spaces and the need for a local inf-sup condition as seen in [46].
@@
In this article, along with the LPS method, we have added another velocity stabilisation to control the normal jump of the solution. This helps to further stabilise the solution. Moreover, we also need pressure stabilisation to stabilise the pressure gradient. In a nutshell, the LPS-HHO method is a combination of a usual HHO method for the Oseen problem combined with the above stabilisations along with an upwind term. Compared to [1], we have proven that the LPS stabilisation term in the formulation helps to prove a stability result under a stronger SUPG-like norm. Moreover, in this article, the presence of normal jump stabilisation in the discrete scheme gives epsilon robust error bounds. This can be seen in inequality (4).
@@
The rest of the article is organised as follows: Section 1 defines the Oseen problem along with some notation and preliminaries. Section 2 deals with the discrete HHO formulation of the Oseen equations. Section 3 provides the proof for the discrete well-posedness of the system in Section 2. Section 5 provides a priori error estimates. The numerical results are provided in Section 6. The notation $a\lesssim b$ means that there exists a generic constant $C$ independent of the meshsize such that $a\leq Cb$. We abbreviate $a\lesssim b\lesssim a$ by $a\sim b$. For $M\subset\Omega$, the $L^{2}$-inner product on $L^{2}(M)$ is denoted by $(\cdot,\cdot)_{M}$ and $L^{2}$-norm by $\|\cdot\|_{M}$. We omit the subscript for the domain specification when $M=\Omega$. We extend the definition $(\cdot,\cdot)_{M}$ naturally to vector-valued functions as the sum of component-wise inner products. The analysis is done on standard $k$th order Sobolev spaces $H^{k}(\Omega)$ with the standard norm $\left\|\cdot\right\|_{k}=\sum_{|\alpha|\leq k}\left\|D^{\alpha}\cdot\right\|_{L^{2}(\Omega)}$. The Sobolev space $H^{1}(\Omega)$ with zero trace is denoted by $H^{1}_{0}(\Omega)$. The space of square-integrable functions with zero mean is denoted by $L^{2}_{0}(\Omega)$. For $M\subset\Omega,k\geq 0$ and $1\leq p\leq\infty$, let .$\|.\|_{k,p,M}$ be the norm on the $k$th order Sobolev space $W^{k,p}(M)$. For $p=2$ we denote the norm by .$\|.\|_{k,M}$.
@@ 1 Continuous Problem, Notations and Preliminaries
In this section, we introduce the weak formulation of the Oseen problem (0.1) and some preliminaries. Let $\boldsymbol{V}=[H^{1}_{0}(\Omega)]^{d}$ be the velocity space and $Q=L^{2}_{0}(\Omega)$ be the pressure space. The weak formulation for the Oseen problem (0.1) is given by: Find $\boldsymbol{u}\in\boldsymbol{V}$ and $p\in Q$ such that
@@
$a(\boldsymbol{u},\boldsymbol{v})$ $:=\epsilon(\nabla\boldsymbol{u},\nabla\boldsymbol{v})+((\boldsymbol{b}\cdot\nabla)\boldsymbol{u},\boldsymbol{v})+(\sigma\boldsymbol{u},\boldsymbol{v}),$
$b(q,\boldsymbol{u})$ $:=(q,{\rm div~{}}\boldsymbol{u}).$
Using the fact that ${\rm div~{}}\boldsymbol{b}=0$ and $\sigma>0$, one can show that the bilinear form $a(\boldsymbol{u},\boldsymbol{v})$ is coercive. It is well known that the bilinear form $b(q,\boldsymbol{u})$ is inf-sup stable for $\boldsymbol{u}\in[H^{1}_{0}(\Omega)]^{d}$ and $q\in L^{2}_{0}(\Omega)$. Therefore, the existence and uniqueness of the problem (1.1) can be shown using the Babuška-Brezzi condition, see [42, Chapter IV]. An equivalent formulation for (1.1) seeks $(\boldsymbol{u},p)\in\boldsymbol{V}\times Q$ such that
@@
Consider a decomposition $\mathcal{T}_{h}$ of the domain $\Omega$ that consists of a finite collection of nonempty disjoint open polyhedral cells $T$ such that $\overline{\Omega}=\cup_{T\in\mathcal{T}_{h}}\overline{T}$ and $h=\max_{T\in\mathcal{T}_{h}}h_{T}$, where $h_{T}$ is the diameter of $T$. We follow [25] for various notation and definitions related to the mesh sequence $\mathcal{T}_{h}$. A closed subset $F$ of $\Omega$ is defined to be a mesh face if it is a subset of an affine hyperplane $H_{F}$ with positive $(d-1)$-dimensional Hausdorff measure and if either of the following two statements holds true: (i) There exist $T_{1}(F)$ and $T_{2}(F)$ in $\mathcal{T}_{h}$ such that $F\subset\partial T_{1}(F)\cap\partial T_{2}(F)\cap H_{F}$; in this case, the face $F$ is called an internal face; (ii) There exists $T(F)\in\mathcal{T}_{h}$ such that $F\subset\partial T(F)\cap\partial\Omega\cap H_{F}$; in this case, the face $F$ is called a boundary face. The set of mesh faces is a partition of the mesh skeleton, that is, $\cup_{T\in\mathcal{T}_{h}}\partial T=\cup_{F\in\mathcal{F}_{h}}\bar{F}$, where $\mathcal{F}_{h}:=\mathcal{F}_{h}^{i}\cup\mathcal{F}_{h}^{b}$ is the collection of all faces, which is the union of the set of all internal faces $\mathcal{F}_{h}^{i}$ and the set of all boundary faces $\mathcal{F}_{h}^{b}$. Let $h_{F}$ denote the diameter of $F\in\mathcal{F}_{h}$. For each $T\in\mathcal{T}_{h}$, the set $\mathcal{F}_{T}:=\{F\in\mathcal{F}_{h}\,|\,F\subset\partial T\}$ denotes the collection of all faces contained in $\partial T$. Following [30, Definition 1], we assume that the mesh sequence $(\mathcal{T}_{h})_{h>0}$ is admissible in the sense that, for all $h>0$, $\mathcal{T}_{h}$ admits a matching simplicial submesh $\mathfrak{T}_{h}$ (i.e., every cell and face of $\mathfrak{T}_{h}$ is a subset of a cell and a face of $\mathcal{T}_{h}$, respectively) so that the mesh sequence $(\mathcal{T}_{h})_{h>0}$ is shape-regular in the usual sense and all the cells and faces of $\mathcal{T}_{h}$ have a uniformly comparable diameter to the cell and face of $\mathcal{T}_{h}$ to which they belong. Owing to [28, Lemma 1.42], for $T\in\mathcal{T}_{h}$ and $F\in\mathcal{F}_{T}$, $h_{F}$ is comparable to $h_{T}$ in the sense that
$\varrho^{2}h_{T}\leq h_{F}\leq h_{T},$
where $\varrho$ is the mesh regularity parameter. Moreover, there exists an integer $N_{\partial}$ depending on $\varrho$ and $d$ such that (see [28, Lemma 1.41])
$\underset{T\in\mathcal{T}_{h}}{\text{max}}\,\text{card}(\mathcal{F}_{T})\leq N_{\partial}.$
Next, we define the hybrid discrete spaces on the decomposition $\mathcal{T}_{h}$. For any bounded domain $S$, let $P^{k}(S)$ denote the space of polynomials defined on $S$ of degree at most $k\geq 0$. Let $P^{k}(\mathcal{T}_{h})$ denote the piecewise $k$ degree polynomial functions defined on $\Omega$. The local degrees of freedom on each polygon $T\in\mathcal{T}_{h}$ is given by
$\underline{\boldsymbol{U}}_{T}^{k}:=\{\underline{\boldsymbol{v}}_{T}=(\boldsymbol{v}_{T},(\boldsymbol{v}_{F})_{F\in\mathcal{F}_{T}}):\boldsymbol{v}_{T}\in[P^{k}(T)]^{d}\text{ and }\boldsymbol{v}_{F}\in[P^{k}(F)]^{d},F\in\mathcal{F}_{T}\}.$
@@
The restriction of $\underline{\boldsymbol{v}}_{h}\in\underline{\boldsymbol{U}}_{h}^{k}$ on a polygon $T$ is denoted by $\underline{\boldsymbol{v}}_{T}$. For $\underline{\boldsymbol{v}}_{h}\in\underline{\boldsymbol{U}}_{h}^{k}$, the broken polynomial function $\boldsymbol{v}_{h}\in[P^{k}(\mathcal{T}_{h})]^{d}$ is defined as $\boldsymbol{v}_{h}|_{T}:=\boldsymbol{v}_{T}$. The local interpolation operator $\underline{\boldsymbol{I}}_{T}^{k}:[H^{1}(T)]^{d}\rightarrow\underline{\boldsymbol{U}}_{T}^{k}$ is given by
@@
where $\pi_{T}^{k}$ and $\pi_{F}^{k}$ are $L^{2}$ orthogonal projection onto $[P^{k}(T)]^{d}$ and $[P^{k}(F)]^{d}$ respectively. Similarly, the global interpolation operator $\underline{\boldsymbol{I}}_{h}^{k}\boldsymbol{v}:[H^{1}(\Omega)]^{d}\rightarrow\underline{\boldsymbol{U}}_{h}^{k}$ is defined as follows:
@@
We recall some standard inequalities that will be used throughout the article.
Inverse Inequality: [25, Lemma 1.28] There exists a positive constant $C$ independent of the meshsize $h_{T}$ such that for any $v_{h}\in P^{k}(T)$ we have
@@
Trace Inequality:  [28, pp. 27] There exists a positive constant $C$ independent of the meshsize $h_{T}$ such that
@@
Discrete Poincaré inequality: [25, Lemma 2.15] There exists a positive constant $C$ independent of $h_{T}$ such that for $\underline{\boldsymbol{v}}_{h}\in\underline{\boldsymbol{U}}_{h,0}^{k}$ we have
$\left\|\boldsymbol{v}_{h}\right\|\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}\left\|\nabla\boldsymbol{v}_{T}\right\|_{T}^{2}+\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\frac{1}{h_{F}}\|\boldsymbol{v}_{F}-\boldsymbol{v}_{T}\|_{F}^{2}\Big{)}^{1/2}.$ (1.3)
Approximation property of $L^{2}$ orthogonal projection:  [28, lemma 1.58] The $L^{2}$-projection $\pi_{T}^{k}$ satisfies the following approximation property: for any $\boldsymbol{v}\in H^{s}(T)$ with $s\in\{1,2,...,k+1\}$
@@
Let the outward unit normal component for a polygon $T\in\mathcal{T}_{h}$ be denoted by $\boldsymbol{n}_{T}$. Similarly, the outward unit normal for a face $F\in\mathcal{F}_{T}$ is given by $\boldsymbol{n}_{TF}$ such that $\boldsymbol{n}_{TF}:=\boldsymbol{n}_{T}\rvert_{F}$. Moreover, the normal component of the convection term on a face $F\in\mathcal{F}_{T}$ is defined as $\boldsymbol{b}_{TF}:=\boldsymbol{b}\rvert_{F}\cdot\boldsymbol{n}_{TF}$. The jump of a scalar-valued function $v$ on a face $F$ shared by two polygons $T_{1}$ and $T_{2}$ is given by
@@
The sign of $\left[\hskip-3.5pt\left[v\right]\hskip-3.5pt\right]$ is adjusted according to the direction of the outward normal. For a vector-valued functions $\boldsymbol{u}=(u_{1},\ldots,u_{d})$ and $\boldsymbol{v}=(v_{1},\ldots,v_{d})$
$\nabla\boldsymbol{v}:=\begin{pmatrix}\frac{\partial v_{i}}{\partial x_{j}}\end{pmatrix}_{d\times d},\hskip 14.22636pt\frac{\partial\boldsymbol{v}}{\partial\boldsymbol{n}}:=(\nabla\boldsymbol{v})\boldsymbol{n}\;\;{\rm and}\;\;(\nabla\boldsymbol{u},\nabla\boldsymbol{v}):=\int_{\Omega}\sum_{i,j=1}^{d}\frac{\partial u_{i}}{\partial x_{j}}\frac{\partial v_{i}}{\partial x_{j}}{\rm~{}dx}.$
For a domain $S$, we define $(\boldsymbol{u},\boldsymbol{v})_{S}:=\int_{S}\boldsymbol{u}\cdot\boldsymbol{v}{\rm~{}dx}$. Let $|v|$ denote the modulus function of $v$. The positive and negative part of $v$ is defined as